# The Grand Challenge of Quantum Applications

### v2

@@ Introduction
Period finding [Shor1999Polynomial] Factoring and discrete log [Shor1999Polynomial] Exponential Classically
@@
QAOA [Farhi2014Quantum] QAOA for 8-SAT [Boulebnane2024Solving] Super-quadratic Classically
Physics simulation Out-of-time-order correlators [Google2025] Exponential Quantumly
Table 2: The short list of Stage II results that are currently known to the authors of this perspective for which there is an exponential-size ensemble with an average-case super-quadratic quantum speedup. This is in contrast to problems such as FeMoco or the Hubbard model where there is effectively a finite number of distinct calculations we expect to have quantum advantage. We only include Stage II results that are classically or quantumly verifiable (and hence, exclude results such as random circuit sampling) as we believe that possessing one of these features is necessary for utility. Also worth mentioning but not included in this table is work on proofs of quantumness based on trapdoor claw-free functions where verification is only efficient using hidden side information [brakerski2021cryptographic, kahanamoku2022classically, brakerski2020simpler, arabadjieva2024single, zhu2023interactive].
@@
We consider a quantum algorithm to have reached Stage II only when one can generate concrete instances of a problem that are expected to exhibit quantum advantage. For example, this would be satisfied if one identifies an ensemble of problem instances that can be efficiently sampled for which a quantum algorithm gives a super-quadratic speedup over the best known classical algorithms, with probability vanishing at most polynomially in the problem size (e.g., they are average-case hard classically). While this is a demanding criterion, several results have met it. Examples of papers that have addressed Stage II include Shor’s algorithm [Shor1999Polynomial], the DQI algorithm applied to optimal polynomial intersection (OPI) [Jordan2025Optimization], algorithms for tensor PCA [Hastings2020Classical] and speedups for other number-theoretic problems like Pell’s equation [Hallgren2007Polynomial]. One could also consider using a quantum computer to efficiently sample hard instances, although we are not aware of any examples of this kind.
@@
A worst-case hardness construction can be considered as a starting gun for the search for a distinctive ensemble of instances on which the quantum algorithm exhibits an average-case speedup. For example, the OTOC problem studied in [Google2025] is BQP-Complete, and there is also a natural distribution of instances that appears to have an average-case speedup. Similarly, there is complexity-theoretic evidence that sampling the QAOA algorithm is classically hard in the worst-case [Farhi2016Quantum] and later studies determined concrete ensembles on which QAOA outperforms classical optimizers [Boulebnane2024Solving].
@@
However, our view is that fixating on asymptotic scaling may be the wrong lens for assessing the practical value of quantum computing in chemistry. Molecules and the correlation lengths of materials are intrinsically finite; thus, an algorithm’s utility depends on its performance on these specific instances, not its behavior at an infinite limit. Furthermore, while some classical methods may have favorable polynomial scaling in theory, they can be completely impractical for high-precision chemistry in practice. For example, the projected entangled pairs state (PEPS) method (the basis for claims that classical methods may efficiently solve electronic structure made in [Lee2023Evaluating]) has never been successfully applied to molecular electronic structure due to its prohibitively high scaling. In three dimensional space, heuristic PEPS contraction scales as roughly as the twelfth power of the bond dimension [Xie2017Optimized] and numerics for the 2D Hubbard model growing in one direction from [Lee2023Evaluating] (which use the simple update method [Corboz2016Variational] that frequently does not converge) indicate that for that system, one needs the bond dimension to grow roughly inverse linearly in the target relative error. In chemistry the scaling is likely even worse due to higher spatial dimensions and often one targets a fixed absolute error (“chemical accuracy”) with molecular energies growing super-linearly.
This brings us to a more pragmatic perspective. For certain high interest systems like FeMoco, we know that despite intense effort, no classical method has solved its electronic structure to the accuracy needed for definitive chemical insight [Berry2025Rapid, Zhai2021Low, Xiang2024Distributed]. We also know that we can efficiently prepare quantum states with a very high overlap on its true ground state [Berry2025Rapid]. Therefore, while the theoretical debate over exponential quantum speedup remains unresolved, having a large-scale quantum computer today would allow us to learn things about the molecule that are currently beyond our reach.
@@
Chemical dynamics (2nd quantization + vibrations) Singlet-fission organic solar cells† [motlagh2025quantum] 4 states 11 modes Evaluated vs. MCTDH [xie2015full] 1100 [motlagh2025quantum] $3\cdot 10^{7}$
@@
Dynamic structure factor evaluation (2nd quantization) Battery redox reaction by Li2MnO3 ($10^{4}$ shots[kunitsa2025quantum] 18 orbitals Plausible beyond RAS methods [fomichev2025fast] 100 [kunitsa2025quantum] $9\cdot 10^{11}$
@@
Chemical dynamics (1st quantization) Reaction of NH3 and BF3 [da2025comprehensive] 40 electrons Plausible but not assessed 800 [da2025comprehensive] $2\cdot 10^{15}$
@@
Table 4: Non-comprehensive compendium of Stage IV research progress in quantum simulation. † indicates estimates are only for single-shot dynamics (thus, more repetitions would be required to estimate an observable). Applications are roughly classified into different time horizons based on the reported number of physical $10^{-3}$ error-rate superconducting qubits (nearest: $\sim 10^{5}$, medium: $\sim 10^{6}$, furthest: $\geq 10^{7}$) needed for a first demonstration of useful quantum advantage on a surface code. One thing to note is that while there are many high value applications requiring the simulation of ab initio solid-state models, such simulations are quite expensive and more research is needed to bring down costs. Where unavailable, we compare across different architectures by making a subjective classification based on the reported number of logical qubits and Toffoli gates (or equivalent), which may not capture routing overheads, opportunities to parallelize non-Clifford operations, or other space-time tradeoffs. Some other resource estimates worth mentioning include Refs. [agrawal2024quantifying, nguyen2024quantum, watts2024fullerene, clinton2024towards].
@@ Acknowledgements
The authors thank the following people for helpful discussions or feedback which helped shape our perspective on these matters: David Gosset, Matt Harrigan, Sid Jain, Stephen Jordan, Robin Kothari, Tony Metger, Hartmut Neven, Matt Reagor, Alexander Schmidhuber, Vadim Smelyanskiy, Rolando Somma, Alec White, Adam Zalcman and Mark Zhandry.

### v3

@@ Introduction
Period finding [Simon1997Power] Factoring and discrete log [Shor1999Polynomial] Exponential Classically
@@
Quant. approx. opt. algo. (QAOA) [Farhi2014Quantum] QAOA for 8-SAT [Boulebnane2024Solving] Super-quadratic Classically
Random circuit sampling (RCS) [Boixo2018Characterizing] RCS out-of-time-order correlators [Google2025] Exponential Quantumly
Table 2: The short list of Stage II results that are currently known to the authors of this perspective for which there is an exponential-size ensemble with an average-case super-quadratic quantum speedup. This is in contrast to problems such as FeMoco or the Hubbard model where there is effectively a finite number of distinct calculations we expect to have quantum advantage. Also worth mentioning but not included in this table is work on proofs of quantumness based on trapdoor claw-free functions where verification is only efficient using hidden side information [brakerski2021cryptographic, kahanamoku2022classically, brakerski2020simpler, arabadjieva2024single, zhu2023interactive].
@@
We consider a quantum algorithm to have reached Stage II only when one can generate concrete instances of a problem that are expected to exhibit verifiable quantum advantage. For example, this would be satisfied if one identifies an ensemble of quantumly verifiable problem instances that can be efficiently sampled for which a quantum algorithm gives a super-quadratic speedup over the best known classical algorithms, with probability vanishing at most polynomially in the problem size (e.g., they are average-case hard classically). While this is a demanding criterion, several results have met it. Examples of papers that have addressed Stage II include Shor’s algorithm [Shor1999Polynomial], the DQI algorithm applied to optimal polynomial intersection (OPI) [Jordan2025Optimization], algorithms for tensor PCA [Hastings2020Classical] and speedups for other number-theoretic problems like Pell’s equation [Hallgren2007Polynomial]. One could also consider using a quantum computer to efficiently generate or sample hard instances, although we are not aware of any examples of this kind.
@@
However, our view is that fixating on asymptotic scaling may be the wrong lens for assessing the practical value of quantum computing in chemistry. Molecules and the correlation lengths of materials are intrinsically finite; thus, an algorithm’s utility depends on its performance on these specific instances, not its behavior at an infinite limit. Furthermore, while some classical methods may have favorable polynomial scaling in theory, they can be completely impractical for high-precision chemistry in practice. For example, the projected entangled pairs state (PEPS) method (the basis for claims that classical methods may efficiently solve strongly correlated electronic structure made in [Lee2023Evaluating]) has never been successfully applied to molecular electronic structure due to its prohibitively high scaling. This brings us to a more pragmatic perspective. For certain high interest systems like FeMoco, we know that despite intense effort, no classical method has solved its electronic structure to the accuracy needed for definitive chemical insight [Berry2025Rapid, Zhai2021Low, Xiang2024Distributed]. We also know that we can efficiently prepare quantum states with a very high overlap on its true ground state [Berry2025Rapid]. Therefore, while the theoretical debate over exponential quantum speedup remains unresolved, having a large-scale quantum computer today would allow us to learn things about the molecule that are currently beyond our reach.
@@
Chemical dynamics (2nd quantization + vibrations) Singlet-fission organic solar cells† [motlagh2025quantum] 6 states 21 modes Evaluated vs. MCTDH [xie2015full] 154 [motlagh2025quantum] $3\cdot 10^{6}$
@@
X-ray absorption spectroscopy (2nd quantization) Structure of Li2MnO3 battery cathode [kunitsa2025quantum] 18 orbitals Plausible beyond RAS methods [fomichev2025fast] 100 [fomichev2025fast] $5\cdot 10^{11}$
@@
Chemical dynamics (1st quantization) Reaction of NH3 and BF3 [da2025comprehensive] 40 electrons Plausible but not assessed 800 [da2025comprehensive] w.o. ancillae $2\cdot 10^{15}$
@@
Table 4: Non-comprehensive compendium of Stage IV research progress in quantum simulation. Dynamics simulation estimates include number of repetitions required to estimate an observable unless marked by †  which indicates that estimates are only for a single shot. Applications are roughly classified into different time horizons based on the reported number of physical $10^{-3}$ error-rate superconducting qubits (nearest: $\sim 10^{5}$, medium: $\sim 10^{6}$, furthest: $\geq 10^{7}$) needed for a first demonstration of useful quantum advantage on a surface code. One thing to note is that while there are many high value applications requiring the simulation of ab initio solid-state models, such simulations are quite expensive and more research is needed to bring down costs. Where unavailable, we compare across different architectures by making a subjective classification based on the reported number of logical qubits and Toffoli gates (or equivalent), which may not capture routing overheads, opportunities to parallelize non-Clifford operations, or other space-time tradeoffs. Some other resource estimates worth mentioning include Refs. [agrawal2024quantifying, nguyen2024quantum, watts2024fullerene, clinton2024towards, malpathak2025trotter, baker2024spindefect, loaiza202infraredspectroscopy, casares2025quantumalgorithmsdetectodmractive].
@@ Acknowledgements
The authors thank the following people for helpful discussions or feedback which helped shape our perspective on these matters: Garnet Chan, David Gosset, Matt Harrigan, Sid Jain, Stephen Jordan, Robin Kothari, Tony Metger, Danial Motlagh, Hartmut Neven, Matt Reagor, Juan Miguel Arrazola, Alexander Schmidhuber, Vadim Smelyanskiy, Rolando Somma, Guifre Vidal, Alec White, Adam Zalcman and Mark Zhandry.